All house sales transactions (in R1000) concluded in two cities during a particular week are
checked. The selling price of each house is noted. The following information is obtained:
Town A Town B
1
Expert's answer
2021-10-12T02:04:17-0400

Assume the missing data is:



Find the test statistic value.

Solution:

"H_0:\mu_1=\mu_2
\\H_1:\mu_1>\mu_2"

Test statistic,"z_0=\dfrac{(\bar x_1-\bar x_2)-(\mu_1-\mu_2)}{\sqrt{\dfrac{\sigma_1^2}{n_1}+\dfrac{\sigma_2^2}{n_2}}}"

"=\dfrac{(13585-13600)-(0)}{\sqrt{\dfrac{1024}{51}+\dfrac{1024}{34}}}
\\=-2.1172"